. NEED ANSWER ASAP / ANSWER NEVER USED BEFORE a.) $46.50 per share is the current...

.

NEED ANSWER ASAP / ANSWER NEVER USED BEFORE

a.)

$46.50 per share is the current price for Foster Farms' stock. The dividend is projected to increase at a constant rate of 5.50% per year. The required rate of return on the stock, rs, is 9.00%. What is the stock's expected price 3 years from today?

Select the correct answer.

a. $54.60
b. $53.96
c. $54.28
d. $54.92

e. $55.24

b.)

Orwell building supplies' last dividend was $1.75. Its dividend growth rate is expected to be constant at 32.00% for 2 years, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the current stock price?

Select the correct answer.

a. $46.00
b. $44.57
c. $47.43
d. $48.86

e. $50.29

c.)

Carby Hardware has an outstanding issue of perpetual preferred stock with an annual dividend of $3.80 per share. If the required return on this preferred stock is 6.5%, at what price should the preferred stock sell?

Select the correct answer.

a. $58.46
b. $57.89
c. $57.32
d. $59.03
e. $56.75
Answer #1

a.P3=Current price*(1+Growth rate)^3

=46.5*(1.055)^3

=$54.60(Approx)..

b.D1=(1.75*1.32)=2.31

D2=(2.31*1.32)=3.0492

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(3.0492*1.06)/(0.12-0.06)

=53.8692

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.31/1.12+3.0492/1.12^2+53.8692/1.12^2

=$47.43(Approx)

c.Current value=Annual dividend/required return

=(3.8/0.065)

=$58.46(Approx).